OpenStudy (anonymous):

Compute each angle of the given triangle. Where necessary, use a calculator and round to one decimal place. a = 31, b = 9, c = 37 ∠A = ? ° ∠B = ? ° ∠C = ? °

OpenStudy (wmj259):

\[c ^{2}=a ^{2}+b ^{2}-2abcos(C)\] 

OpenStudy (wmj259):

This formula I gave you is the Law of Cosines, which lets you to solve for each angle of a triangle if your given (3 side lengths) or (2 side lengths and 1 angle between them)

OpenStudy (anonymous):

I worked it out, but idk how to get the top number

OpenStudy (anonymous):

I know the bottom number is 260

OpenStudy (wmj259):

Angle A?

OpenStudy (anonymous):

Yup the denominator is 260

OpenStudy (anonymous):

Idk how to get the numerator

OpenStudy (wmj259):

So if you rewrite the law of cosines to find the angle your formula would then be \[\frac{ c ^{2}-a ^{2}-b ^{2} }{-2ab}=\cos(C)\] taking the inverse cosine would then give you \[C=\cos^{-1} (\frac{ c ^{2}-a ^{2}-b ^{2} }{-2ab})\]

OpenStudy (anonymous):

\[\cos C=\frac{ a^2+b^2-c^2 }{ 2ab }\]

OpenStudy (wmj259):

@surjithayer, yes thats correct with the -1 from the denominator distributed into the numerator and reorganized.

OpenStudy (anonymous):

If C is obtuse then cos C is negative.

OpenStudy (anonymous):

so, its 233/260 for A

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\cos A=\frac{ b^2+c^2-a^2 }{ 2bc }\]

OpenStudy (anonymous):

How about Cos B?

OpenStudy (wmj259):

that would be cos(B)=(a^2+c^2-b^2)/2ac